A point charge of +3.5 μC is placed 0.800 m to the left of the second, identical point charge.

What is the magnitudes and directions of forces that each charge exerts on the other?


1
Expert's answer
2022-01-12T08:29:07-0500

Two identical point charges will have the equal magnitudes of forces that each charge exerts on the other:


"F_{12}=F_{21}=\dfrac{kq_1q_2}{r^2},""F_{12}=F_{21}=\dfrac{9.0\times10^9\ \dfrac{N\times m^2}{C^2}\times(3.5\times10^{-6}\ C)^2}{(0.8\ m)^2},""F_{12}=F_{21}=0.17\ N."

Since two charges are both positively charged they will repel each other. As a result, the force on charge "q_1" is opposite in direction to the force it exerts on charge "q_2".